Word problems!! Can you help me solve: Find three consecutive integers whose sum divided by the difference between the third number and the first number will be 3 more than the middle number.
Thank you!

Answer by stanbon(75887)

Find three consecutive integers whose sum divided by the difference between the third number and the first number will be 3 more than the middle number.
-----
1st: x-1
2nd: x
3rd: x+1
------------------
Equation:
(3x)/(2) = x+3
---
3x = 2x+6
x = 6
----
1st: x-1 = 5
2nd: x = 6
3rd: x+1 = 7
=================
Check:
(18)/2 = 6+3
9 = 9
===================
Cheers,
Stan H.